# A multiplexed control architecture for superconducting qubits with row-column addressing

### 06 Mar 2024 v1

@@ III.2 Two-qubit gates
Table 2: Two-qubit schemes with two control pulses. Similar to the qubit addressing, note that for two-qubit gates based on frequency-tunable coupler or bus, static DC flux biases (not listed here) are generally needed for biasing the coupler at their idle points. This can be achieved by using an on-chip programmable magnetic memory, as shown in the left inset of Fig. 3(a).
@@
Figure 7: The numerical verification of the two-qubit gate operations based on two-tone flux modulations. In the numerical analysis, we consider that two transmon qubits (labeled by $Q_{a}$ and $Q_{b}$) are coupled via a tunable bus (labeled by $Q_{bus}$), and the system state is denoted by $|Q_{a}Q_{bus}Q_{b}\rangle$. The qubit frequencies are $\omega_{a(b)}/2\pi=5.05(5.25)\,\rm GHz$ and the qubit anharmonicity is $\eta_{a(b)}/2\pi=-250\,\rm MHz$. The maximum frequency and the anharmonicity of the tunable bus are $\omega_{bus}/2\pi=6.20\,\rm GHz$ and $\eta_{bus}/2\pi=-200\,\rm MHz$, respectively. The qubit-bus coupling strength is $g/2\pi=100\,\rm MHz$. The dependence of bus frequency on the two-tone flux modulation is approximated by $\omega_{bus}(\Phi)=(\omega_{bus}-\eta_{bus})\sqrt{|\cos(\pi\Phi(t)/\Phi_{0})|}+\eta_{bus}$ with $\Phi(t)=\Phi_{p}(t)[\sin(\omega_{p1}t)+\sin(\omega_{p2}t+\phi_{0})]$ (for simplicity, here $\phi_{0}=0$) and the modulation frequency $\omega_{p1}/2\pi=2.0\,\rm GHz$. Here, similar to Fig. 4(c), raised cosine flat-top pulses are used and the peak modulation amplitude is $\Phi_{p}$. (a) Population $P_{100}$ (upper panel) and $P_{001}$ (bottom panel) versus the modulation amplitude $\Phi_{p}$ and the modulation frequency $\omega_{p2}$ with the qubit prepared in state $|100\rangle$. The ramp time is 15 ns and the total gate length is 100 ns. Dashed orange lines indicate the available regions for realizing $\sqrt{\textmd{iSWAP}}$ gates while the red star is for iSWAP gates. Here, the two-qubit gate is actuated when the difference frequency of the two modulations matches the qubit-qubit detuning. (b) Population $P_{101}$ (upper panel) versus the modulation amplitude $\Phi_{p}$ and the modulation frequency $\omega_{p2}$ with the system prepared in state $|101\rangle$. The bottom panel shows the associated conditional phase. The ramp time is 25 ns and the total gate length is 200 ns. Here, the CZ gate is realized when the difference frequency of the two modulations matches the detuning between the energy levels of $|101\rangle$ and $|200\rangle$. Note that discontinuities near the horizontal line at $\Phi_{p}/\Phi_{0}=0.13$ result from the parasitic interactions between the energy levels of $|101\rangle$ and $|110\rangle$.
@@
As discussed in Sec. II, to be compatible with the multiplexed control architecture, we consider two-qubit gates based on driving single qubit or a coupler in two-coupled qubit systems. Accordingly, Table 2 summarizes three possible two-qubit gate schemes supporting the row-column addressing:
@@
(DT1) For two fixed-frequency qubits coupled via a coupler (bus), two-qubit gates can be obtained by only tuning the coupler (bus) frequency Collodo2020 ; Xu2020 ; Zhao2022b ; Goto2022 ; Campbell2023 . Given the shared flux pulses, each coupler (bus) can be individually tuned by using the on-chip programmable magnetic memory Johnson2010 ; Bunyk2014 ; Zhao2023 , as shown in Fig. 3(a). Note here that with the help of the local memory, in principle, a single Z line, which is shared by all the couplers or buses, is adequate for performing parallel two-qubit gates.
@@
(DT2) Given the nonlinear dependence of the transmon qubit or transmon coupler frequency on the flux bias, two-qubit gates can be realized by applying two-tone flux modulations on the qubits or the coupler. When the the sum or difference frequency of the two-tone modulations match the subharmonic of the qubit-qubit detuning, two-qubit gates can be actuated. For example, considering two fixed-frequency transmon qubits coupled via tunable bus McKay2016 , Figure 7 shows two-qubit gates, such as iSWAP, $\sqrt{\textmd{iSWAP}}$, and CZ gates can be obtained by the two-tone modulation of the bus frequency. This two-tone modulation scheme can also be applied to other coupler circuits Mundada2019 ; Han2020 ; Kubo2023 . In Appendix B, we further show that two-qubit gates can be obtained by applying the two-tone flux modulation on one of the two coupled transmon qubits.
@@
(DF1) For two-qubit gates based on driving a single qubit, the three-wave mixer could be used to generate the desired effective single-qubit drive. However, considering the existing techniques Frattini2017 ; Chapman2023 ; Liu2014 ; Zhao2018 , the strength of the effective drive (here the strength is about $\sim 1\,\rm MHz$) is too weak to support the implementation of a successful, microwave-activated two-qubit gate, such as a cross-resonance (CR) gate Chow2011 . Therefore, here new physical components should be introduced for generating effective drives with large amplitudes.
@@
Given the above discussions, the scheme of (DT2) is more feasible than the other two schemes while the scheme of (DF1) is the most challenging one as new physical components should be demonstrated. Besides the three discussed schemes, similar to the sing-qubit gate scheme of (SF2), two-qubit gates, which are actuated by applying two-tone microwave drives to a qubit or a coupler, should be potential solutions for supporting the row-column addressing and are thus worth exploring in future works.
@@ IV Challenges towards large-scale multiplexed control architectures
In state-of-the-art superconducting quantum processors, several issues, such as distortions of control signals Gustavsson2013 ; Barends2014 , crosstalk between control signals Acharya2023a , and temporal fluctuations of qubit parameters and coherence times (due to, e.g., two-level systems) Muller2019 , are not yet to be well addressed for achieving reliable, accurate quantum computing. This becomes even more challenging within the multiplexed control architecture.
To be more specific, currently, independent control allows various active approaches to be employed for alleviating the detrimental effects of these issues on qubits. On the contrary, due to the control limitations, these approaches cannot be directly applied to the multiplexed control architecture. Thus, rather than attempting to calibrate away these issues, instead, one might turn to fully suppress or eliminate these issues at the qubit chip level. Nevertheless, given state-of-the-art technologies, these progresses are unlikely to be achieved soon. Additionally, at the control pulse level, robust quantum control could be explored to provide resilience against these issues Dong2010 , while at the circuit level, gate compilations with adaptations to these issues can be developed Wang2022 ; Wang2022a .
Besides the well-recognized challenges above, there exist two challenges that are particularly important in achieving high-fidelity qubit control with the multiplexed control but are rarely involved in the traditional independent control architecture. In the following, we will discuss the two challenges to be faced when scaling up the multiplexed control architecture.
@@ IV.1 The non-uniformity of qubit parameters
As illustrated in Sec. II and Sec. III, here the gate pulses are shared by multiple qubits, thus the gate condition for each qubit is distinctively intertwined with each other. For clarity, we assume that the coupling efficiencies between control lines and qubits are all the same and focus on the single-qubit gates (similar results can also be obtained for the two-qubit gates discussed in Sec. III.2). Thus, for the single-qubit gate scheme illustrated in Sec. III.1, given fixed-length drives with the same pulse shape, gates can only be actuated when the gate conditions
@@
are satisfied. Here $\Omega_{i}$ ($\Omega_{i}$) and $\omega_{i}$ ($\omega_{i}$) denote the amplitude and frequency of the two drive pulses applied to the qubit, which is located at the intersection of the $i$th row line and the $j$th column line. As mentioned in Sec. III.1, in the following, we focus on $\sqrt{\textmd{X}}$ gates, which provide more flexibility in setting pulse parameters (similarly, for two-qubt gates, one might prefer $\sqrt{\textmd{CZ}}$ gates Shi2023 or $\sqrt{\textmd{iSWAP}}$ gates, see, e.g., Fig. 7).
Given the condition of Eq. (1) and fixed drive amplitudes, pulse solutions for the scheme, which relies on the resonance condition $f(\omega_{i},\omega_{j})=\omega_{ij}$ (see, e.g., Fig. 8 in Appendix A), such as (SF2), (ST1), and (ST2), can in principle exist. This is because in a square qubit lattice comprising $n\times n$ qubits, $2n$ different frequencies should be adequate for actuating parallel gate operations. However, even if solutions exist, considering that the resonance conditions are generally intertwined with the drive amplitudes and the non-uniformity of qubit parameters, such as the coupling efficiency between qubits and control lines (i.e., leading to the non-uniformity of the drive amplitude), are ubiquitous in reality, whether and how such solutions can be realized practically in large-scale systems should be open questions and should be an issue worthy of further exploration in further works. In the following, we also consider an alternative approach, i.e., reducing the number of conditions, allowing us to explore another extreme situation.
As shown in Fig. 2(a), in 2D square qubit lattices, the two sets of shared row-column lines allow us to selectively address neighboring qubits, i.e., data qubits and syndrome qubits. In this way, the qubit frequency allocation in each diagonal can follow a zigzag pattern, e.g., date qubits at a frequency band with a typical value of $\omega_{D}$ and syndrome qubits at one another frequency band with a typical value of $\omega_{S}$. Thus, the gate conditions for data qubits and syndrome qubits are decoupled to each other and we can deal with data qubits and syndrome qubits separately. Accordingly, in the following, we focus on the data qubits, and the gate conditions for date qubits can be approximately reduced to a single one
@@
where the subscript $D$ denotes the gate condition for data qubits. Finding the solution of the above equation is equivalent to control of an ensemble of qubits, where the qubit frequencies could be different from qubit to qubit (given state-of-the-art technologies Hertzberg2021 ; Pappas2024 , the non-uniformity of qubit frequency can be suppressed to the level of $10\,\rm MHz$ Hertzberg2021 ), using uniform control pulses, i.e., the parameters of the pulse in each row or column are the same. However, in fact, besides the qubit frequency and the coupling efficiency, the non-uniformity of qubit parameters also results from other factors, such as the qubit anharmonicity and the phase difference among control lines. Therefore, considering all these non-uniformity of qubit parameters, whether one can find the pulse solution of Eq. (2) for multiplexed control of large-scale qubit systems highly hinges on the uniformity of qubit parameters.
As shown in Fig. 3(a), in principle, on-chip programmable magnetic memory can be used to mitigate most of the above-discussed non-uniform issues Johnson2010 ; Zhao2023 but its compatibility with high-fidelity qubit control has not yet been demonstrated. Meanwhile, robust quantum control could also be explored to optimize the control pulse against the remaining non-uniformity issues Dong2010 ; Shao2024 . Thus, given state-of-the-art quantum technologies and the above discussions, we expect that only taking one of the two types of gate solutions, which supports the gate conditions in Eqs. (1) and (2), alone is challenging to enable high-fidelity multiplexed qubit control at scale. One might choose to take the combination of such two solutions, thus allowing the adaptations to various issues, such as the non-uniformity issue.
@@ IV.2 Gate calibration
While individual gate calibrations and benchmarking can still be possible as in the independent control architecture Kelly2018 ; Klimov2020 , the multiplexed control should lead to a significant degradation of the efficiency and the performance that can be reached. This is because that the gate calibration procedures for each qubit are intertwined with each other. Alternatively, one might also consider fine-tuning up gates at the quantum circuit level, such as stabilize measurement circuits Kelly2016 , where the detection events can be used for informing and guiding the optimization of gate parameters.
@@
In conclusion, we introduce a multiplexed control architecture for superconducting qubits with shared row-column lines, which could in principle provide an efficient approach for parallel controlling $N$ qubits with $O(\sqrt{N})$ control lines. We also propose various single- and two-qubit gate schemes that are both compatible with the row-column qubit addressing and the existing superconducting qubit technologies. Given the inherent parallelism of this architecture, we show that the multiplexed control is suitable for the implementation of structured quantum circuits. As an immediate application, we show that the architecture can be specifically tailored to execute the quantum error correction with surface code. We envision that a small-scale demonstration of the multiplexed qubit control could be feasible with the current technologies and hope that our work could motivate further experimental and theoretical research in incorporating shared control into scalable quantum information processing with superconducting qubits.
@@ Appendix A Nonlinear dependence on the frequency difference in the two-tone drive scheme
Figure 8: Nonlinear dependence on the frequency difference of the two-tone microwave drive. The left (right) panel shows the population $P_{0}$ (the leakage $L_{1}$) versus the drive-qubit detunings $\delta_{1}$ and $\delta_{2}$ (white regions are where leakage below $10^{-5}$). As in Fig. 3, the qubit frequency and the anharmonicity are $\omega_{q}/2\pi=5.25\,\rm GHz$ and $\eta_{q}/2\pi=-250\,\rm MHz$, respectively. Here, the peak drive parameters are $\Omega_{1(2)}/2\pi=25\rm MHz$. Dashed grey lines indicate the available regions for realizing $\sqrt{\textmd{X}}$ gates while the red star is for X gates
For the two-tone microwave drives studied in Fig. 3(d), Figure 8 shows the population $P_{0}$ and the leakage $L_{1}$ versus the drive-qubit detunings $\delta_{1}$ and $\delta_{2}$.
@@
Figure 9: The two-qubit gate operations based on applying two-tone flux modulations to qubits. We consider that two transmon qubits (labeled by $Q_{a}$ and $Q_{b}$) are coupled directly and a two-tone flux modulation is applied to the frequency-tunable qubit $Q_{a}$. The system state is denoted by $|Q_{a}Q_{b}\rangle$. Here, as in Fig. 7(a), the raised cosine flat-top pulses are used with a total gate length of 100 ns and a ramp time of 15 ns. The left (right) panel shows population $P_{10}$ ($P_{01}$) versus the modulation amplitude $\Phi_{p}$ and the modulation frequency $\omega_{p2}$ with the system prepared in state $|10\rangle$. Dashed orange lines indicate the available regions for realizing $\sqrt{\textmd{iSWAP}}$ gates while the red star is for iSWAP gates.
Here, we consider that one frequency-tunable transmon qubit (labeled by $Q_{a}$) is coupled fixedly to a fixed-frequency transmon qubit (labeled by $Q_{b}$) with the coupling strength of $g/2\pi=5.5\,\rm MHz$. The qubit frequencies and the qubit anharmonicities are $\omega_{a(b)}/2\pi=5.05(5.25)\,\rm GHz$ and $\eta_{a(b)}/2\pi-250\,\rm MHz$, respectively. When applying the two-tone flux modulation to $Q_{a}$, the qubit frequency is approximated by $\omega_{a}(\Phi)=(\omega_{a}-\eta_{a})\sqrt{|\cos(\pi\Phi(t)/\Phi_{0})|}+\eta_{a}$ with $\Phi(t)=\Phi_{p}(t)[\sin(\omega_{p1}t)+\sin(\omega_{p2}t+\phi_{0})]$ Here, the modulation frequency $\omega_{p1}/2\pi=50\,\rm MHz$ and $\phi_{0}=0$ for simplicity. Figure 9 shows that under such a two-tone flux modulation, two-qubit gates, such as $\sqrt{\textmd{iSWAP}}$ gates and iSWAP gates, can be actuated.

### 20 Mar 2024 v2

@@ III.2 Two-qubit gates
Table 2: Two-qubit schemes with two control pulses. Similar to the qubit addressing, note that for two-qubit gates based on frequency-tunable coupler or bus, static dc flux biases (not listed here) are generally needed for biasing the coupler or bus at their idle points. This can be achieved by using an on-chip programmable magnetic memory, as shown in the left inset of Fig. 3(a).
@@
As discussed in Sec. II, to be compatible with the multiplexed control architecture, we consider two-qubit gates based on driving single qubit or a coupler in two-coupled qubit systems. Table 2 summarizes three possible two-qubit gate schemes supporting the row-column addressing:
@@
(DT1) For two fixed-frequency qubits coupled via a coupler (bus), two-qubit gates can be obtained by only tuning the coupler (bus) frequency Collodo2020 ; Xu2020 ; Zhao2022b ; Goto2022 ; Campbell2023 . Given the shared flux pulses, each coupler (bus) can be individually tuned by using the on-chip programmable magnetic memory Johnson2010 ; Bunyk2014 ; Zhao2023 , as shown in Fig. 3(a). Note that with the help of local memory, in principle, only Z lines shared by the couplers or buses are needed to perform parallel two-qubit gates.
@@
(DT2) By leveraging the nonlinear dependence of the transmon qubit or transmon coupler frequency on the flux bias, two-qubit gates can be realized by applying two-tone flux modulations on the qubits or the coupler. When the the sum or difference frequency of the two modulations match the subharmonic of the qubit-qubit detuning, two-qubit gates can be actuated. For example, considering two fixed-frequency transmon qubits coupled via tunable bus McKay2016 , Figure 7 shows that two-qubit gates, such as iSWAP, $\sqrt{\textmd{iSWAP}}$, and CZ gates can be obtained by the two-tone modulation of the bus frequency. This two-tone modulation scheme is adaptable to other coupler circuits Mundada2019 ; Han2020 ; Kubo2023 . In Appendix B, we further show that two-qubit gates can be obtained by applying the two-tone flux modulation to one of the coupled transmon qubits.
@@
(DF1) For two-qubit gates based on driving a single qubit, the three-wave mixer could be used to generate the desired effective single-qubit drive. However, considering the existing techniques Frattini2017 ; Chapman2023 ; Liu2014 ; Zhao2018 , the strength of the effective drive (here the strength is about $\sim 1\,\rm MHz$) is too weak to support the successful implementation of microwave-activated two-qubit gates, such as cross-resonance (CR) gates Chow2011 . Therefore, new physical components that enable the three-wave mixing process should be introduced to generate effective drives with large amplitudes.
@@
Given the above discussions, the scheme of (DT2) emerges as the most viable option among the three schemes, whereas the scheme of (DF1) poses the greatest challenge, necessitating the demonstration of new physical components. In addition to the three discussed schemes, similar to the sing-qubit gate scheme of (SF2), two-qubit gates, which are actuated by applying two-tone microwave drives to a qubit or a coupler, should be potential solutions for supporting the row-column addressing and are thus worth exploring in future works.
@@ IV Challenges towards large-scale multiplexed control architectures
In state-of-the-art superconducting quantum processors, several issues, such as distortions in control signals Gustavsson2013 ; Barends2014 , crosstalk among control signals Acharya2023a , and temporal fluctuations of qubit parameters and coherence times (due to factors like two-level systems) Muller2019 , are not yet to be well addressed for achieving reliable, accurate quantum computing. This becomes even more challenging within the multiplexed control architecture.
To be more specific, currently, independent control allows various active approaches to be employed for alleviating the detrimental effects of these issues on qubits. However, due to the control limitations, these approaches cannot be directly applied to address similar issues within the multiplexed control architecture. Consequently, rather than merely attempting to calibrate away these issues, instead, one might turn to fully suppress or eliminate these issues at the qubit chip level. Nevertheless, given state-of-the-art technologies, significant advancements in this regard is unlikely to be achieved soon. Additionally, at the level of control pulses, robust quantum control could be explored to provide resilience against these issues Dong2010 , while at the circuit level, gate compilations with adaptations to address these issues can be developed Wang2022 ; Wang2022a .
In addition to the aforementioned well-recognized hurdles, there exist two new challenges that are particularly important in achieving high-fidelity qubit control within the multiplexed control architecture but are rarely involved in the traditional independent control architecture. In the following, we will discuss the two challenges to be faced when considering scaling up the multiplexed control architecture.
@@ IV.1 The non-uniformity of qubit parameters
As illustrated in Sec. II and Sec. III, here the gate pulses are shared by multiple qubits, thus the gate condition for each qubit is distinctively intertwined with each other. For clarity, we assume that the coupling efficiencies between control lines and qubits are all the same and focus on the single-qubit gates (similar results can also be obtained for the two-qubit gates discussed in Sec. III.2). In the single-qubit gate schemes illustrated in Sec. III.1, with fixed-length drives featuring the same pulse shape, gates can only be actuated when the gate conditions
@@
are satisfied. Here $\Omega_{i}$ ($\Omega_{i}$) and $\omega_{i}$ ($\omega_{i}$) denote the amplitude and frequency of the two drive pulses applied to the qubit (with the frequency of $\omega_{ij}$), which is located at the intersection of the $i$th row line and the $j$th column line. As mentioned in Sec. III.1, in the following, we focus on $\sqrt{\textmd{X}}$ gates, which provide more flexibility in setting pulse parameters (similarly, for two-qubt gates, one might prefer $\sqrt{\textmd{CZ}}$ gates Shi2023 or $\sqrt{\textmd{iSWAP}}$ gates, see, e.g., Fig. 7).
Under the condition of Eq. (1) and given fixed drive amplitudes, pulse solutions for the scheme, which relies on the resonance condition $f(\omega_{i},\omega_{j})=\omega_{ij}$ (see, e.g., Fig. 8 in Appendix A), such as (SF2), (ST1), and (ST2), can in principle exist. This is because in a square qubit lattice comprising $n\times n$ qubits, $2n$ distinct frequencies should suffice for actuating parallel gate operations. However, even if solutions exist, considering that the resonance conditions are generally intertwined with the drive amplitudes and the non-uniformity of qubit parameters, such as the coupling efficiency between qubits and control lines (i.e., leading to the non-uniformity of the drive amplitude), are ubiquitous in reality, whether and how such solutions can be realized practically in large-scale systems should be open questions. This issue warrants further exploration in further works. In the following, we also consider an alternative approach, i.e., reducing the number of conditions, allowing us to explore another extreme (’trivial’) situation.
In the context of 2D square qubit lattices depicted in Fig. 2(a), the utilization of two sets of shared row-column lines allow us to selectively address neighboring qubits, i.e., data qubits and syndrome qubits. In this way, the qubit frequency allocation in each diagonal can follow a zigzag pattern, e.g., where date qubits at a frequency band with a typical value of $\omega_{D}$ and syndrome qubits at a separate frequency band with a typical value of $\omega_{S}$. Consequently, the gate conditions for data qubits and syndrome qubits are decoupled to each other, facilitating a separate treatment for each type of qubits. Hence, in the following, we focus on the data qubits with gate conditions potentially being reduced to a single criterion denoted by
@@
where the subscript $D$ indicates the gate condition for data qubits. Finding the solution of the above equation is equivalent to control of an ensemble of qubits, where the qubit frequencies could be different from qubit to qubit (given state-of-the-art technologies Hertzberg2021 ; Pappas2024 , the non-uniformity of qubit frequency can be suppressed to the level of $10\,\rm MHz$ Hertzberg2021 ), using uniform control pulses, i.e., the parameters of the pulse in each row or column are the same.
However, in fact, besides the qubit frequency and the coupling efficiency, the non-uniformity of qubit parameters also results from other factors, such as the qubit anharmonicity and the phase difference among control lines. Therefore, considering all these non-uniformity of qubit parameters, whether one can find the pulse solution of Eq. (2) for multiplexed control of large-scale qubit systems crucially hinges on the uniformity of qubit parameters. Furthermore, if these non-uniformity issues can be effectively addressed, the multiplexed control can be simplified to a trivial situation, eliminating the need for row-column addressing and two-pulse control configurations.
As shown in Fig. 3(a), in principle, on-chip programmable magnetic memory can be used to mitigate most of the above-discussed non-uniform issues Johnson2010 ; Zhao2023 but its compatibility with high-fidelity qubit control has not yet been demonstrated. Meanwhile, robust quantum control could also be explored to optimize the control pulse against these non-uniformity issues Dong2010 ; Hai2022 ; Shao2024 . Given state-of-the-art quantum technologies and the above discussions, we expect that solely adopting one of the two types of gate solutions supporting the gate conditions in Eqs. (1) and (2) presents a formidable hurdle in achieving high-fidelity multiplexed qubit control at scale. Therefore, one might choose to take the combination of such two solutions, thus enabling the adaptations to address various issues, including the non-uniformity issue.
@@ IV.2 Gate calibration
While individual gate calibrations and benchmarking may still be possible as in the independent control architecture Kelly2018 ; Klimov2020 , the multiplexed control should lead to a significant degradation of the efficiency and the performance that can be reached. This is because that the gate calibration procedures for each qubit are intertwined with each other. Alternatively, one might also consider fine-tuning up gates at the quantum circuit level, such as stabilize measurement circuits Kelly2016 , where the detection events can be used for informing and guiding the optimization of gate parameters.
Moreover, when calibrating gate parameters tailored to implement distinct gates on individual qubits, additional overhead is required. As illustrated in Fig. 2(a), given the $\sqrt{\textmd{X}}$ gates, arbitrary single-qubit gates, including identity operations, can be directly optimized by tuning up the Z gates. However, as discussed in Sec. II, selectively activating two-qubit gates on qubit subgroups, such as implementing standard two-qubit gates (e.g., iSWAP-type gates and CZ gates) on selective qubits while applying identity operations to the inactive qubits, presents a nontrivial task for gate tune-up.
These challenges generally stem from the tradeoff between line overhead and control flexibility. In order to compensate for the degradation in control flexibility, additional gates or gate patterns need to be tuned up, potentially leading to an increase in circuit depths Tan2024 .
@@
In conclusion, we introduce a multiplexed control architecture for superconducting qubits with shared row-column lines. This architecture could in principle provide an efficient approach for parallel controlling $N$ qubits with $O(\sqrt{N})$ control lines. We also propose various single- and two-qubit gate schemes that are both compatible with the row-column qubit addressing scheme and the existing superconducting qubit technologies. Leveraging the inherent parallelism of this architecture, we show that the multiplexed control is suitable for the implementation of structured quantum circuits. As an immediate application, we show that the architecture can be specifically tailored to execute the quantum error correction with surface code.
We envision that a proof-of-concept demonstration of multiplexed qubit control on a small scale could be feasible with the current technologies and hope that our work could motivate further experimental and theoretical research in incorporating shared control into scalable quantum information processing with superconducting qubits.
@@ Acknowledgements.
Figure 8: Nonlinear dependence on the frequency difference of the two-tone microwave drive. The left (right) panel shows the population $P_{0}$ (the leakage $L_{1}$) versus the drive-qubit detunings $\delta_{1}$ and $\delta_{2}$ (white regions are where leakage below $10^{-5}$). Here, the peak drive parameters are $\Omega_{1(2)}/2\pi=25\rm MHz$. Dashed grey lines indicate the feasible regions for realizing $\sqrt{\textmd{X}}$ gates while the red stars indicate the gate parameters for implementing X gates.
@@ Appendix A Nonlinear dependence on the frequency difference in the two-tone drive scheme
For the two-tone microwave drives of (SF2) studied in Fig. 3(d), Figure 8 shows the population $P_{0}$ and the leakage $L_{1}$ versus the drive-qubit detunings $\delta_{1}$ and $\delta_{2}$, demonstrating the nonlinear dependence of the single-qubit dynamics on the frequency difference of the two drives. Here, as in Fig. 3, in the numerical analysis, the qubit frequency and the anharmonicity are $\omega_{q}/2\pi=5.25\,\rm GHz$ and $\eta_{q}/2\pi=-250\,\rm MHz$, respectively.
@@
Here, we consider that one frequency-tunable transmon qubit, labeled by $Q_{a}$, is coupled fixedly to a fixed-frequency transmon qubit, labeled by $Q_{b}$, with the coupling strength of $g/2\pi=5.5\,\rm MHz$. The qubit frequencies and the qubit anharmonicities are $\omega_{a(b)}/2\pi=5.05(5.25)\,\rm GHz$ and $\eta_{a(b)}/2\pi-250\,\rm MHz$, respectively. When applying the two-tone flux modulation to $Q_{a}$, the qubit frequency is approximated by $\omega_{a}(\Phi)=(\omega_{a}-\eta_{a})\sqrt{|\cos(\pi\Phi(t)/\Phi_{0})|}+\eta_{a}$ with $\Phi(t)=\Phi_{p}(t)[\sin(\omega_{p1}t)+\sin(\omega_{p2}t+\phi_{0})]$. Here, the modulation frequency is $\omega_{p1}/2\pi=50\,\rm MHz$ and $\phi_{0}=0$ for simplicity. As in Fig. 7(a), the raised cosine flat-top pulses are used, with a total gate length of 100 ns and a ramp time of 15 ns. Figure 9 demonstrates that under such a two-tone flux modulation, two-qubit gates, such as $\sqrt{\textmd{iSWAP}}$ gates and iSWAP gates, can be actuated.
Figure 9: The two-qubit gate operations based on applying two-tone flux modulations to qubits. We consider that two transmon qubits, labeled by $Q_{a}$ and $Q_{b}$, are coupled directly and a two-tone flux modulation is applied to the frequency-tunable qubit $Q_{a}$. The system state is denoted by $|Q_{a}Q_{b}\rangle$. The left (right) panel shows population $P_{10}$ ($P_{01}$) versus the modulation amplitude $\Phi_{p}$ and the modulation frequency $\omega_{p2}$ with the system prepared in state $|10\rangle$. Dashed orange lines indicate the available regions for realizing $\sqrt{\textmd{iSWAP}}$ gates while the red star is for iSWAP gates.